A mass of 60kg supported by a chain 13m long, is pulled a perpendicular distance of 5m from the original line of support. Calculate the horizontal force required?


1
Expert's answer
2020-12-29T15:31:55-0500

Explanations & Calculations

  • Assume the mass is hung by the chain of 13m & pulled perpendicularly.
  • Considering the equilibrium of the system at a distance of 5m from the starting point, the force needed can be calculated.
  • In the beginning, the needed force is zero as there is no force acting opposite to what we apply, and gradually a component of its own weight starts acting as an opposite force to what we apply as it moves in a circular arch from the start to the final position.



  • At the equilibrium,

Applying Newton's second for the verticle equilibrium "\uparrow" ,

"\qquad\qquad\qquad
\begin{aligned}
\small T\sin\theta-mg&=\small 0\\
\small T\sin \theta &= \small 60kg\times9.8=588N\cdots(1) 
\end{aligned}"

From the horizontal equilibrium,

"\qquad\qquad\qquad
\begin{aligned}
\small T\cos\theta &= \small F\cdots(2)
\end{aligned}"

  • By (1)/(2),

"\qquad\qquad\qquad
\begin{aligned}
\small \tan\theta&= \small \frac{588}{F}\\
\small F&= \small \frac{588}{\tan\theta}
\end{aligned}"

  • By simple geometry at the equilibrium position, "\tan\theta" can be found. Chain is 13m, horizontal is 5m hence the verticle side 12m (by pythagoras's)
  • Thereofre,

"\qquad\qquad\qquad
\begin{aligned}
\small F&= \small \frac{588}{\frac{12}{5}}=245N
\end{aligned}"